The acceleration of a cylinder plough and the displacement is b=m√x where m is constant. The velocity is in the direction of the displacement and the velocity and displacement are both zero when time is zero. calculate displacement, velocity and acceleration as a function of time.
1
Expert's answer
2021-10-24T01:17:27-0400
"\frac{d^2x}{dt^2}=m\sqrt{x}\\x(t)=0\\v(t)=0"